# Vision-based Learning for Drones: A Survey

### 08 Dec 2023 v1

@@ Abstract
Drones as advanced cyber-physical systems are undergoing a transformative shift with the advent of vision-based learning, a field that is rapidly gaining prominence due to its profound impact on drone autonomy and functionality. Different from existing task-Specific surveys, this review offers a comprehensive overview of vision-based learning in drones, emphasizing its pivotal role in enhancing their operational capabilities. We start by elucidating the fundamental principles of vision-based learning, highlighting how it significantly improves drones’ visual perception and decision-making processes. We then categorize vision-based control methods into indirect, semi-direct, and end-to-end approaches from the perception-control perspective. We further explore various applications of vision-based drones with learning capabilities, ranging from single-agent systems to more complex multi-agent and heterogeneous system scenarios, and underscore the challenges and innovations characterizing each area. Finally, we explore open questions and potential solutions, paving the way for ongoing research and development in this dynamic and rapidly evolving field. With growing large language models (LLMs) and embodied intelligence, vision-based learning for drones provides a promising but challenging road towards artificial general intelligence (AGI) in 3D physical world.
{IEEEImpStatement}
The rapid development of vision-based learning in drones represents a significant leap forward in the realm of autonomous systems, with applications extending across various sectors. This comprehensive survey offers an in-depth look at how vision-based learning enhances drone autonomy, decision-making, and operational capacity. With the first attempt to emphasize integration with large language models and embodied intelligence, the review navigates through the achievements and challenges towards artificial general intelligence in the physical world. This work not only underscores the technological advancements but also addresses the open questions and possible solutions for vision-based drones, setting a foundation for future innovations in drone autonomy and its application in complex, real-world environments.
@@
Drones; Intelligent Systems; Robotics Learning
@@ 1 Introduction
  • The autonomy of drones. Drone autonomy achieves autonomous navigation and task execution. It requires real-time perception and onboard computation capabilities. Visual sensors and efficient online planning with learning have extended drone autonomy and intelligence [13, 14], even beating world-level champions in drone racing [15].
Among these directions, vision-based drones with learning capabilities are attracting more insights (see the growing trend in Fig. 2) and emerging towards artificial general intelligence (AGI) in the 3D physical world, especially when integrated with rapid-growing large language models (LLMs) [16] and embodied intelligence [17]. Existing vision-based drone-related surveys focus only on specific tasks and applications, such as UAV navigation [18, 19], vision-based UAV landing [20], obstacle avoidance [21], vision-based inspection with UAVs [22, 23], and autonomous drone racing [24], which limits the understanding of vision-based learning for drones from a general perspective. Therefore, this survey paper provides a comprehensive review of vision-based learning for drones to deliver a more general view of current drone autonomy technologies, including background, visual perception, vision-based control, applications and challenges, and open questions with potential solutions. To summarize, our main contributions are:
@@
  • We discussed the current state of vision-based control methods for drones and categorized them into indirect, semi-direct, and end-to-end methods from the perception-action perspective. This perspective helps to understand vision-based control methods and differentiate them with better features.
@@
Organization: The rest of this survey is organized as follows: Section 2 discusses the concept of vision-based learning for drones and the core components; Section 3 summarizes object detection with visual perception and its application to vision-based drones; We introduce the vision-based control methods for drones and categorize them in Section 4; The applications and challenges of vision-based learning for drones are discussed in Section 5; We list the open questions faced by vision-based learning for drones and potential solutions in Section 6; Section 7 summarizes and concludes this survey.
@@ 2.1 Vision-based Learning Drones
Currently, vision-based drones with learning capabilities, which utilize cheaper and more flexible visual sensors and efficient learning algorithms, have achieved remarkable high-level performance in a series of standardized visual perception and decision-making tasks, such as agile flight control [25], navigation [26] and obstacle avoidance [27]. For instance, using only depth cameras, inertial measurement units (IMU) and a lightweight onboard computer, the vision-based drone in [25] succeeded in performing high-speed flight in unseen and unstructured wild environments. The controller of the vision-based drone in [25] was trained from a high-fidelity simulation and transferred to a physical platform. Following that, the Swift system was developed in [15] to achieve world champion-level autonomous drone racing with a tracking camera and a shallow neural network. Such kinds of vision-based agile drones are leading the future of drones due to their perception and learning capabilities in complex environments. Similarly, with event cameras and a trained narrow neural network-EVDodgeNet, the vision-based drone in [28] was able to dodge multiple dynamic obstacles (balls) during flight. Afterwards, to improve the perception onboard in the real world, an uncertainty estimation module was trained with the Ajna network [29], which significantly increased the generalization capability of learning-based control for drones. These cases showcase the power of learning algorithms in improving the agility and perception capabilities of vision-based drones.
A typical vision-based drone consists of 3 parts (see Fig. 3): (1) Visual perception: sensing the environment around the drone via monocular cameras or stereo cameras; (2) Image processing: extracting features from an observed image sequence and output specific patterns or information, such as navigation information, depth information, object information; (3) Flight controller: generating high-level and low-level commands for drones to perform assigned missions. Image processing and flight controller are generally conducted on the onboard computer, while visual perception relies on the performance of visual sensors.
Figure 3: General framework of vision-based drones
@@ 2.2.1 LIght Detection And Ranging (LIDAR)
LIDAR is a kind of active range sensor that relies on the calculated time of flight (TOF) between the transmitted and received beams (laser) to estimate the distance between the robot and the reflected surface of objects [30]. Based on the scanning mechanism, LIDAR can be divided into solid-state LIDAR, which has a fixed field of view without moving parts, and surrounding LIDAR, which spins to provide a 360-degree horizontal view. Surrounding LIDAR is also referred to as “laser scanning” or “3D scanning”, which creates a 3D representation of the explored environment using eye-safe laser beams. A typical LIDAR (see Fig. 4) consists of laser emitters, laser receivers, and a spinning motor. The vertical field of view (FOV) of a LIDAR is determined by the number of vertical arrays of lasers. For instance, a vertical array of 16 lasers scanning 30 degrees gives a vertical resolution of 2 degrees in a typical configuration.
LIDAR has recently been used on drones for mapping [31, 32, 33], power grid inspection [34], pose estimation [35] and object detection [36, 37]. LIDAR provides sufficient and accurate depth information for drones to navigate in cluttered environments. However, it is bulky and power-hungry and does not fit within the payload restrictions of agile autonomous drones. Meanwhile, using raycast representation in a simulation environment makes it hard to match the inputs of a real LIDAR device, which brings many challenges for the Sim2Real transfer when a learning approach is considered.
@@ 2.2.2 Camera
Compared to LIDAR, cameras provide a cheap and lightweight way for drones to perceive the environment. Cameras are external passive sensors used to monitor the drone’s geometric and dynamic relationship to its task, environment or the objects that it is handling. Cameras are commonly used perception sensors for drones to sense environment information, such as objects’ position and a point cloud map of the environment. In contrast to the motion capture system, which can only broadcast global geometric and dynamic pose information within a limited space from an offboard synchronized system, cameras enable a drone to fly without space constraints. Cameras can provide positioning for drone navigation in GPS-denied environments via visual odometry (VO) [38, 39, 40] and visual simultaneous localization and mapping systems (V-SLAM) [41, 42, 43]. Meanwhile, object detection and depth estimation can be performed with cameras to obtain the relative positions and sizes of obstacles. However, to avoid dynamic obstacles, even physical attacks, the agile vision-based drone poses fundamental challenges to visual perception. Motion blur, sparse texture environments, and unbalanced lighting conditions can cause the loss of feature detection in the VIO and object detection. LIDAR and event cameras [44] can partially address these challenges. However, LIDAR and event cameras are either too bulky or too expensive for agile drone applications. Considering the agility requirement of physical attack avoidance, lightweight dual-fisheye cameras are used for visual perception. With dual fisheye cameras, the drone can achieve better navigation capability [45] and omnidirectional visual perception [46]. Some sensor fusion and state estimation techniques are required to alleviate the accuracy loss brought by the motion blur.
@@ 3 Object Detection with Visual Perception
Object detection is a pivotal module in vision-based drones when handling complex missions such as inspection, avoidance, and search and rescue. Object detection is to find out all the objects of interest in the image and determine their position and size. Object detection is one of the core problems in the field of computer vision (CV). Nowadays, the applications of object detection include face detection, pedestrian detection, vehicle detection, and terrain detection in remote sensing images. Object detection has always been one of the most challenging problems in the field of CV due to the different appearances, shapes, and poses of various objects, as well as the interference of factors such as illumination and occlusion during imaging. At present, the object detection algorithm can be roughly divided into two categories: multi-stage (two-stage) algorithm, whose idea is to first generate candidate regions and then perform classification, and one-stage algorithm, the idea of which is to directly apply the algorithm to the input image and output the categories and corresponding positions. Beyond that, to retrieve 3D positions, depth estimation has been a popular research subbranch related to object detection whether using monocular [58] or stereo depth estimation [59]. For a very long time, the core neural network module (backbone) of object detection has been the convolutional neural network (CNN). CNN is a classic neural network in image processing that originates from the study of the human optic nerve system. The main idea is to convolve the image with the convolution kernel to obtain a series of reorganization features, and these reorganization features represent the important information of the image. In this way, CNN not only has the ability to recognize the image but also effectively decreases the requirement for computing resources. Recently, vision transformers (ViTs) [60], originally proposed for image classification tasks, have been extended to the realm of object detection [61]. These models demonstrate superior performance by utilizing the self-attention mechanism, which processes visual information non-locally. However, a major limitation of ViTs is their high computational demand. This presents difficulties in achieving real-time inference, particularly on platforms with limited resources like drones.
@@ 3.1 Multi-stage Algorithms
Faster R-CNN is an improvement on the basis of Fast R-CNN. In order to solve the problem of SSA, the SSA that generates RP in Fast R-CNN is replaced by a Region Proposal Network (RPN) and uses a model that integrates RP generation, feature extraction, object classification and object box regression. RPN is a fully convolutional network that simultaneously predicts object boundaries at each location. RPN is trained end-to-end to generate high-quality region proposals, which are then detected by Fast R-CNN. At the same time, RPN and Fast R-CNN share convolutional features. Meanwhile, in the feature extraction stage, Faster R-CNN uses a convolutional neural network. The model achieves $73.2\%$ and $70.4\%$ mean Average Precision (mAP) per category on the PASCAL VOC 2007 and 2012 datasets, respectively. Faster R-CNN has been greatly improved in speed than Fast R-CNN, and the accuracy has reached the SOTA, and it also fully developed an end-to-end object detection framework. However, Faster R-CNN still cannot achieve real-time object detection. Besides, after obtaining RP, it requires heavy computation for each RP classification.
@@ 3.3 Vision Transformer
When applying object detection algorithms to drone applications, it is necessary to find the best balance between computation speed and accuracy. Besides, massive drone datasets are required for training and testing. Zheng Ye et al. [74] collected an air-to-air drone dataset “Det-Fly” (see Fig. 11) and evaluated air-to-air object detection of a micro-UAV with eight different object detection algorithms, namely RetinaNet [75], SSD, Faster R-CNN, YOLOv3 [76], FPN [77], Cascade R-CNN[78] and Grid R-CNN [79]. The evaluation results in [74] showed that the overall performance of Cascade R-CNN and Grid R-CNN is superior compared to the others. However, the YOLOv3 provides the fastest inference speed among others. Wei Xun et al. [80] conducted another investigation into drone detection, employing the YOLOv3 architecture and deploying the model on the NVIDIA Jetson TX2 platform. They collected a dataset comprising 1435 images featuring various UAVs, including drones, hexacopters, and quadcopters. Utilizing custom-trained weights, the YOLOv3 model demonstrated proficiency in drone detection within images. However, the deployment of this trained model faced constraints due to the limited computational capacity of the Jetson TX2, which posed challenges for effective real-time application.
@@ 4 Vision-based Control
Vision-based control for robotics has been widely studied in recent years, whether for ground robots or aerial robotics such as drones. For drones flying in a GPS-denied environment, visual inertial odometry (VIO) [38, 39, 40] and visual simultaneous localization and mapping systems (SLAM) [41, 42, 43] have been preferred choices for navigation. Meanwhile, in a clustered environment, research on obstacle avoidance [25, 82, 83, 13] based on visual perception has attracted much attention in the past few years. Obstacle avoidance has been a main task for the vision-based control of drones. However, there are few works proposed for vision-based drones flying against adversarial environments, such as physical attacks from birds or intelligent adversaries. It requires omnidirectional visual perception, real-time multi-object detection and tracking, and agile avoidance capabilities for drones to fly away from these adversarial agents.
Obstacle avoidance methods are currently only developed and deployed on drones for static [84, 13, 25, 85] or simple dynamic environments [81, 86, 27], where the paths of the moving obstacles can be easily projected (e.g., moving gates [86] or balls being thrown [27], see Fig. 12) while in flight. As a result, these studies focus solely on non-intelligent and non-perceptive agents and obstacles. Since the upcoming motion sequence can be predicted, the potential area of conflict is restricted, and the drone is only required to make the corresponding avoidance maneuvering decision at the predicted position in time.

### 02 Jan 2024 v2

@@ Abstract
Drones as advanced cyber-physical systems are undergoing a transformative shift with the advent of vision-based learning, a field that is rapidly gaining prominence due to its profound impact on drone autonomy and functionality. Different from existing task-specific surveys, this review offers a comprehensive overview of vision-based learning in drones, emphasizing its pivotal role in enhancing their operational capabilities under various scenarios. We start by elucidating the fundamental principles of vision-based learning, highlighting how it significantly improves drones’ visual perception and decision-making processes. We then categorize vision-based control methods into indirect, semi-direct, and end-to-end approaches from the perception-control perspective. We further explore various applications of vision-based drones with learning capabilities, ranging from single-agent systems to more complex multi-agent and heterogeneous system scenarios, and underscore the challenges and innovations characterizing each area. Finally, we explore open questions and potential solutions, paving the way for ongoing research and development in this dynamic and rapidly evolving field. With growing large language models (LLMs) and embodied intelligence, vision-based learning for drones provides a promising but challenging road towards artificial general intelligence (AGI) in 3D physical world.
@@
Drones, learning systems, robotics learning, embodied intelligence.
@@ 1 Introduction
  • The autonomy of drones. Drone autonomy achieves autonomous navigation and task execution. It requires real-time perception and onboard computation capabilities. Drones that integrate visual sensors, efficient online planning, and learning-based decision-making algorithms have notably enhanced autonomy and intelligence [13, 14], even beating world-level champions in drone racing with a vision-based learning system called Swift [15].
Recently, to improve drone autonomy, vision-based learning drones, which combine advanced sensing with learning capabilities, are attracting more insights (see the rapid growth trend in Fig. 2). With such capabilities, drones are even emerging towards artificial general intelligence (AGI) in the 3D physical world, especially when integrated with rapid-growing large language models (LLMs) [16] and embodied intelligence [17]. Existing vision-based drone-related surveys focus only on specific tasks and applications, such as UAV navigation [18, 19], vision-based UAV landing [20], obstacle avoidance [21], vision-based inspection with UAVs [22, 23], and autonomous drone racing [24], which limits the understanding of vision-based learning in drones from a holistic perspective. Therefore, this survey provides a comprehensive review of vision-based learning for drones to deliver a more general view of current drone autonomy technologies, including background, visual perception, vision-based control, applications and challenges, and open questions with potential solutions. To summarize, our main contributions are:
@@
  • We discussed the current state of vision-based control methods for drones and categorized them into indirect, semi-direct, and end-to-end methods from the perception-control perspective. This perspective helps to understand vision-based control methods and differentiate them with better features.
@@
Organization: The rest of this survey is organized as follows: Section 2 discusses the concept of vision-based learning drones and their core components; Section 3 summarizes object detection with visual perception and its application to vision-based drones; We introduce the vision-based control methods for drones and categorize them in Section 4; The applications and challenges of vision-based learning for drones are discussed in Section 5; We list the open questions faced by vision-based learning drones and potential solutions in Section 6; Section 7 summarizes and concludes this survey.
Figure 3: General framework of vision-based drones
@@ 2.1 Vision-based Learning Drones
A typical vision-based drone consists of three parts (see Fig. 3): (1) Visual perception: sensing the environment around the drone via monocular cameras or stereo cameras; (2) Image processing: extracting features from an observed image sequence and output specific patterns or information, such as navigation information, depth information, object information; (3) Flight controller: generating high-level and low-level commands for drones to perform assigned missions. Image processing and flight controllers are generally conducted on the onboard computer, while visual perception relies on the performance of visual sensors. Vision-based drones have been widely used in traditional missions such as environmental exploration [25], navigation [13], and obstacle avoidance [26, 27]. With efficient image processing and simple path planners, they can avoid dynamic obstacles effectively (see Fig. 4).
Currently, vision-based learning drones, which utilize visual sensors and efficient learning algorithms, have achieved remarkable advanced performance in a series of standardized visual perception and decision-making tasks, such as agile flight control [28], navigation [29] and obstacle avoidance [27]. Various cases have showcased the power of learning algorithms in improving the agility and perception capabilities of vision-based drones. For instance, using only depth cameras, inertial measurement units (IMU) and a lightweight onboard computer, the vision-based learning drone in [28] succeeded in performing high-speed flight in unseen and unstructured wild environments. The controller of the drone in [28] was trained from a high-fidelity simulation and transferred to a physical platform. Following that, the Swift system was developed in [15] to achieve world champion-level autonomous drone racing with a tracking camera and a shallow neural network. Such kinds of vision-based learning drones are leading the future of drones due to their perception and learning capabilities in complex environments. Similarly, with event cameras and a trained narrow neural network-EVDodgeNet, the vision-based learning drone in [30] was able to dodge multiple dynamic obstacles (balls) during flight. Afterwards, to improve the perception onboard in the real world, an uncertainty estimation module was trained with the Ajna network [31], which significantly increased the generalization capability of learning-based control for drones. The power of deep learning in handling uncertain information frees traditional approaches from complex computation with necessary accurate modeling.
@@ 2.2.1 LIght Detection And Ranging (LIDAR)
LIDAR is a kind of active range sensor that relies on the calculated time of flight (TOF) between the transmitted and received beams (laser) to estimate the distance between the robot and the reflected surface of objects [32]. Based on the scanning mechanism, LIDAR can be divided into solid-state LIDAR, which has a fixed field of view without moving parts, and surrounding LIDAR, which spins to provide a 360-degree horizontal view. Surrounding LIDAR is also referred to as “laser scanning” or “3D scanning”, which creates a 3D representation of the explored environment using eye-safe laser beams. A typical LIDAR (see Fig. 5) consists of laser emitters, laser receivers, and a spinning motor. The vertical field of view (FOV) of a LIDAR is determined by the number of vertical arrays of lasers. For instance, a vertical array of 16 lasers scanning 30 degrees gives a vertical resolution of 2 degrees in a typical configuration. LIDAR has recently been used on drones for mapping [33], power grid inspection [34], pose estimation [35] and object detection [36]. LIDAR provides sufficient and accurate depth information for drones to navigate in cluttered environments. However, it is bulky and power-hungry and does not fit within the payload restrictions of agile autonomous drones. Meanwhile, using raycast representation in a simulation environment makes it hard to match the inputs of a real LIDAR device, which brings many challenges for the Sim2Real transfer when a learning approach is considered.
@@ 2.2.2 Camera
Compared to LIDAR, cameras provide a cheap and lightweight way for drones to perceive the environment. Cameras are external passive sensors used to monitor the drone’s geometric and dynamic relationship to its task, environment or the objects that it is handling. Cameras are commonly used perception sensors for drones to sense environment information, such as objects’ position and a point cloud map of the environment. In contrast to the motion capture system, which can only broadcast global geometric and dynamic pose information within a limited space from an offboard synchronized system, cameras enable a drone to fly without space constraints. Cameras can provide positioning for drone navigation in GPS-denied environments via visual odometry (VO) [37, 38, 39] and visual simultaneous localization and mapping systems (V-SLAM) [40, 41, 42]. Meanwhile, object detection and depth estimation can be performed with cameras to obtain the relative positions and sizes of obstacles. However, to avoid dynamic obstacles, even physical attacks like bird chasing, the agile vision-based drone poses fundamental challenges to visual perception. Motion blur, sparse texture environments, and unbalanced lighting conditions can cause the loss of feature detection in the VIO and object detection. LIDAR and event cameras [43] can partially address these challenges. However, LIDAR and event cameras are either too bulky or too expensive for agile drone applications. Considering the agility requirement of physical attack avoidance, lightweight dual-fisheye cameras are used for visual perception. With dual fisheye cameras, the drone can achieve better navigation capability [44] and omnidirectional visual perception [45]. Some sensor fusion and state estimation techniques are required to alleviate the accuracy loss brought by the motion blur.
@@ 3 Object Detection with Visual Perception
Object detection is a pivotal module in vision-based learning drones when handling complex missions such as inspection, avoidance, and search and rescue. Object detection is to find out all the objects of interest in the image and determine their position and size [57]. Object detection is one of the core problems in the field of computer vision (CV). Nowadays, the applications of object detection include face detection, pedestrian detection, vehicle detection, and terrain detection in remote sensing images. Object detection has always been one of the most challenging problems in the field of CV due to the different appearances, shapes, and poses of various objects, as well as the interference of factors such as illumination and occlusion during imaging. At present, the object detection algorithm can be roughly divided into two categories: multi-stage (two-stage) algorithm, whose idea is to first generate candidate regions and then perform classification, and one-stage algorithm, the idea of which is to directly apply the algorithm to the input image and output the categories and corresponding positions. Beyond that, to retrieve 3D positions, depth estimation has been a popular research subbranch related to object detection whether using monocular [58] or stereo depth estimation [59]. For a very long time, the core neural network module (backbone) of object detection has been the convolutional neural network (CNN) [60]. CNN is a classic neural network in image processing that originates from the study of the human optic nerve system. The main idea is to convolve the image with the convolution kernel to obtain a series of reorganization features, and these reorganization features represent the important information of the image. As such, CNN not only has the ability to recognize the image but also effectively decreases the requirement for computing resources. Recently, vision transformers (ViTs) [61], originally proposed for image classification tasks, have been extended to the realm of object detection [62]. These models demonstrate superior performance by utilizing the self-attention mechanism, which processes visual information non-locally [63]. However, a major limitation of ViTs is their high computational demand. This presents difficulties in achieving real-time inference, particularly on platforms with limited resources like drones.
@@ 3.1 Multi-stage Algorithms
Faster R-CNN is an improvement on the basis of Fast R-CNN. In order to solve the problem of SSA, the SSA that generates RP in Fast R-CNN is replaced by a Region Proposal Network (RPN) and uses a model that integrates RP generation, feature extraction, object classification and object box regression. RPN is a fully convolutional network that simultaneously predicts object boundaries at each location. RPN is trained end-to-end to generate high-quality region proposals, which are then detected by Fast R-CNN. At the same time, RPN and Fast R-CNN share convolutional features. Meanwhile, in the feature extraction stage, Faster R-CNN uses a convolutional neural network. The model achieves $73.2\%$ and $70.4\%$ mean Average Precision (mAP) per category on the PASCAL VOC 2007 and 2012 datasets, respectively. Faster R-CNN has been greatly improved in speed than Fast R-CNN, and the accuracy has reached the state-of-the-art (SOTA), and it also fully developed an end-to-end object detection framework. However, Faster R-CNN still cannot achieve real-time object detection. Besides, after obtaining RP, it requires heavy computation for each RP classification.
@@ 3.3 Vision Transformer
When applying object detection algorithms to drone applications, it is necessary to find the best balance between computation speed and accuracy. Besides, massive drone datasets are required for training and testing. Zheng Ye et al. [76] collected an air-to-air drone dataset “Det-Fly” (see Fig. 12) and evaluated air-to-air object detection of a micro-UAV with eight different object detection algorithms, namely RetinaNet [77], SSD, Faster R-CNN, YOLOv3 [78], FPN [79], Cascade R-CNN[80] and Grid R-CNN [81]. The evaluation results in [76] showed that the overall performance of Cascade R-CNN and Grid R-CNN is superior compared to the others. However, the YOLOv3 provides the fastest inference speed among others. Wei Xun et al. [82] conducted another investigation into drone detection, employing the YOLOv3 architecture and deploying the model on the NVIDIA Jetson TX2 platform. They collected a dataset comprising 1435 images featuring various UAVs, including drones, hexacopters, and quadcopters. Utilizing custom-trained weights, the YOLOv3 model demonstrated proficiency in drone detection within images. However, the deployment of this trained model faced constraints due to the limited computation capacity of the Jetson TX2, which posed challenges for effective real-time application.
@@ 4 Vision-based Control
Vision-based control for robotics has been widely studied in recent years, whether for ground robots or aerial robotics such as drones. For drones flying in a GPS-denied environment, visual inertial odometry (VIO) [37, 38, 39] and visual simultaneous localization and mapping systems (SLAM) [40, 41, 42] have been preferred choices for navigation. Meanwhile, in a clustered environment, research on obstacle avoidance [28, 83, 84, 13] based on visual perception has attracted much attention in the past few years. Obstacle avoidance has been a main task for vision-based control as well as for the current learning algorithms of drones.
4.1 Indirect Methods
Indirect methods [85, 27, 86, 87, 88, 89, 90, 13, 91, 92, 93] refer to extracting features from images or videos to generate visual odometry, depth maps, and 3D point cloud maps for drones to perform path planning based on traditional optimization algorithms (see Fig. 14). Obstacle states, such as 3D shape, position, and velocity, are detected and mapped before a maneuver is taken. Once online maps are built or obstacles are located, the drone can generate a feasible path or take actions to avoid obstacles.
SOTA indirect methods generally divide the mission into several subtasks, namely perception, mapping and planning. On the perception side, depth images are always required to generate corresponding distance and position information for navigation. A depth image is a grey-level or color image that can represent the distance between the surfaces of objects from the viewpoint of the agent. Fig. 15 shows color images and corresponding depth images from a drone’s viewpoint. The illuminance is proportional to the distance from the camera. A lighter color denotes a nearer surface, and darker areas mean further surfaces. A depth map provides the necessary distance information for drones to make decisions to avoid static and dynamic obstacles. Currently, off-the-shelf RGB-D cameras, such as the Intel RealSense depth camera D415, the ZED 2 stereo camera, and the Structure Core depth camera, are widely used for drone applications. Therefore, traditional obstacle avoidance methods can treat depth information as a direct input. However, for omnidirectional perception in wide-view scenarios, efficient onboard monocular depth estimation is always required, which is a challenge to address with existing methods.
On the mapping side, a point cloud map [94] or Octmap [95], representing a set of data points in a 3D space is commonly generated. Each point has its own Cartesian coordinates and can be used to represent a 3D shape or an object. A 3D point cloud map is not from the view of a drone but constructs a global 3D map that provides global environmental information for a drone to fly. The point-cloud map can be generated from a LIDAR scanner or many overlapped images combined with depth information. An illustration of an original scene and a point cloud map are shown in Fig. 14, where the drone can travel around without colliding with static obstacles.
Planning is a basic requirement for a vision-based drone to avoid obstacles. Within the indirect methods, planning can be further divided into two categories: one is offline methods based on high-resolution maps and pre-known position information, such as Dijkstra’s algorithm [96], A-star [97], RRT-connect[98] and sequential convex optimization [99]; the other is online methods based on real-time visual perception and decision-making. Online methods can be further categorized into online path planning [90, 13, 91] and artificial potential field (APF) methods [27, 100].